# Scaling Towards the Information Boundary of Instruction Set: InfinityInstruct-Subject Technical Report

### v1

@@ 5.3 Scaling Phenomenon in Label Connectivity Distribution
Figure 15: The connection degree of fine-grained tags with other tags exhibits a Power-Law distribution: Freq(degree=d)${}\sim\text{d}^{\gamma}$.
@@
The presence of this topology highlights the interdependent structure of knowledge in instruction data, which should be considered within the data construction, instruction selection, and model training process. This also highlights the necessity of understanding the distribution and interrelation of instruction contents (Zhao et al., 2025), and the necessity of a high-quality tagging system, which is the foundation of understanding such knowledge structure.
Moreover, prior work (e.g., (Tan et al., 2024)) hypothesizes that the scaling law between model performance and dataset size is originated from the distribution of “latent skills” within the data, if such “latent skills” follow a power-law distribution. Our findings provide empirical support for this hypothesis, offering new insights into the origin of scaling laws between data distribution and model performance. We leave further exploration of this direction to future work, which further emphasizes the importance of the tagging system introduced in this paper.
@@ 6.1 Instruction Data Synthesis
Manually constructed datasets rely on experts to write instructions and responses, such as LIMA (Zhou et al., 2023) and Dolly (Ding et al., 2023). While these datasets are high in quality, they are expensive to scale due to the high cost of human annotation.
@@
Semi-automatic approaches, like Self-Instruct(Wang et al., 2023), Alpaca(Taori et al., 2023), and Evol-Instruct(Muennighoff et al., 2023), expand a small set of human-labeled data using prompt engineering. While they improve scalability, the reliance on handcrafted prompts limits diversity and complexity. Fully automatic methods (e.g., WebInstruct(Yue et al., 2024), backtranslation) extract instruction-like data from web documents with minimal human input, but often lack precise control over coverage and difficulty. Recent studies explore improvements from three angles:
@@
  • Seed Selection and High-Information Filtering: The choice of seed data largely determines the downstream diversity and quality. Traditional methods often use a few high-frequency, simple instructions, leading to limited coverage(Wang et al., 2023). Recent work addresses this by leveraging hierarchical tag systems and multi-dimensional metrics to select high-information seeds—i.e., rare, diverse, and complex instructions that better represent undercovered regions of the task space(Tu et al., 2024; He et al., 2025; Xie et al., 2023). This ensures that synthesized data can expand both the coverage (domain/task diversity) and depth (instruction complexity) of the dataset.
@@
  • Evolution-Based Instruction Generation: To address the need for more challenging instructions, evolution-based methods iteratively expand seed data using strategies like genetic algorithms (Xu et al., 2023) or model feedback loops(Gu et al., 2025). These approaches increase instruction complexity and reasoning depth, while dynamically adjusting generation based on model performance. They also help avoid redundancy by diversifying synthesis paths.
@@
  • Instruction Synthesis Strategies: Most existing synthesis methods rely on prompting large language models with handcrafted or auto-generated templates(Wang et al., 2023; Xu et al., 2023). While effective, this can lead to templated and repetitive outputs. Magpie(Xu et al., 2024) proposes prompt-free instruction synthesis via autoregressive alignment, which improves fluency and semantic richness. Similarly, Web Reconstruction (Jiang et al., 2025) leverages real-world text to generate instruction-style data with higher realism and contextual diversity.
@@
Despite these advances, many datasets still underrepresent complex, multi-step, or domain-specific instructions. Few existing works explicitly address the coverage and depth of instruction distribution in a unified manner. This limits model generalization in scenarios requiring deep reasoning, long-range context, or adaptation to unfamiliar domains (Yu et al., 2024; Zhang et al., 2025). Moreover, while some studies explore feedback-driven data augmentation or generative instruction synthesis (Le et al., 2022), most lack a closed-loop framework that jointly optimizes coverage, difficulty, and model adaptation. The interplay between seed selection, model feedback, and evolution remains underexplored.
@@
To address these gaps, we propose a unified instruction data construction framework that systematically expands both coverage and complexity. By integrating hierarchical labeling, high-information seed filtering, evolution-based synthesis, and deficiency-driven augmentation, our approach enables structured, scalable, and adaptive instruction data generation, laying the foundation for robust and generalizable instruction-tuned models.
@@ 6.2 Model Self-Improvement
Model self-improvement aims to iteratively enhance model capabilities through self-generated data or feedback signals. Representative approaches include self-refinement (Madaan et al., 2023; Zhou et al., 2024), multi-round generation and evaluation loops, and performance-driven data augmentation (Yang et al., 2025). These methods enable models to identify and address their own weaknesses by generating more challenging training samples, thereby facilitating continual improvement.
To ensure data quality, recent studies introduce deficiency diagnosis mechanisms (Lighterness et al., 2024), which analyze model performance on downstream tasks to detect knowledge gaps or skill deficiencies. These insights guide the targeted synthesis of training data(Miller et al., 2025). By incorporating such feedback into the training loop, models and datasets can co-evolve—improving both model performance and data coverage over time(Bauer et al., 2024).
Beyond simply increasing instruction quantity, recent research emphasizes instruction coverage (task/domain diversity) and depth (reasoning complexity). Several works have explored combining hierarchical tagging with complexity control to synthesize increasingly challenging and diverse instructions via evolutionary algorithms (Lu et al., ; Zhao et al., 2025). These strategies enhance the cognitive depth of training data and promote models’ abilities in complex reasoning and compositional generalization.
@@ Appendix A Appendix
Given a dialogue between a user and a conversational assistant, and a classification standard that describes different abilities, determine which one or more major categories from this standard the assistant would need to utilize in order to complete the given dialogue. Note: only list the names of the abilities, and separate each different ability with angle brackets <>absent<>< >. Classification Standard: standard

### v2

@@ 5.3 Scaling Phenomenon in Label Connectivity Distribution
[…]
@@
The presence of this topology highlights the interdependent structure of knowledge in instruction data, which should be considered within the data construction, instruction selection, and model training process. This also highlights the necessity of understanding the distribution and interrelation of instruction contents (zhao2025beyond), and the necessity of a high-quality tagging system, which is the foundation of understanding such knowledge structure.
Moreover, prior work (e.g., (tan2024information)) hypothesizes that the scaling law between model performance and dataset size is originated from the distribution of “latent skills” within the data, if such “latent skills” follow a power-law distribution. Our findings provide empirical support for this hypothesis, offering new insights into the origin of scaling laws between data distribution and model performance. We leave further exploration of this direction to future work, which further emphasizes the importance of the tagging system introduced in this paper.
@@ 6.1 Instruction Data Synthesis
Manually constructed datasets rely on experts to write instructions and responses, such as LIMA (lima2023) and Dolly (ding2023enhancing). While these datasets are high in quality, they are expensive to scale due to the high cost of human annotation.
@@
Semi-automatic approaches, like Self-Instruct(wang2023self), Alpaca(taori2023stanford), and Evol-Instruct(muennighoff2023octopack), expand a small set of human-labeled data using prompt engineering. While they improve scalability, the reliance on handcrafted prompts limits diversity and complexity. Fully automatic methods (e.g., WebInstruct(yue2024mammoth2), backtranslation) extract instruction-like data from web documents with minimal human input, but often lack precise control over coverage and difficulty. Recent studies explore improvements from three angles:
@@
  • Seed Selection and High-Information Filtering: The choice of seed data largely determines the downstream diversity and quality. Traditional methods often use a few high-frequency, simple instructions, leading to limited coverage(wang2023self). Recent work addresses this by leveraging hierarchical tag systems and multi-dimensional metrics to select high-information seeds—i.e., rare, diverse, and complex instructions that better represent undercovered regions of the task space(tu2024resofilter; he2025fine; xie2023data). This ensures that synthesized data can expand both the coverage (domain/task diversity) and depth (instruction complexity) of the dataset.
@@
  • Evolution-Based Instruction Generation: To address the need for more challenging instructions, evolution-based methods iteratively expand seed data using strategies like genetic algorithms (xu2023wizardlm) or model feedback loops(gu2025llm). These approaches increase instruction complexity and reasoning depth, while dynamically adjusting generation based on model performance. They also help avoid redundancy by diversifying synthesis paths.
@@
  • Instruction Synthesis Strategies: Most existing synthesis methods rely on prompting large language models with handcrafted or auto-generated templates(wang2023self; xu2023wizardlm). While effective, this can lead to templated and repetitive outputs. Magpie(xu2024magpie) proposes prompt-free instruction synthesis via autoregressive alignment, which improves fluency and semantic richness. Similarly, Web Reconstruction (jiang2025instruction) leverages real-world text to generate instruction-style data with higher realism and contextual diversity.
@@
Despite these advances, many datasets still underrepresent complex, multi-step, or domain-specific instructions. Few existing works explicitly address the coverage and depth of instruction distribution in a unified manner. This limits model generalization in scenarios requiring deep reasoning, long-range context, or adaptation to unfamiliar domains (yu2024survey; zhang2025out). Moreover, while some studies explore feedback-driven data augmentation or generative instruction synthesis (le2022coderl), most lack a closed-loop framework that jointly optimizes coverage, difficulty, and model adaptation. The interplay between seed selection, model feedback, and evolution remains underexplored.
@@
To address these gaps, we propose a unified instruction data construction framework that systematically expands both coverage and complexity. By integrating hierarchical tagging, high-information seed filtering, evolution-based synthesis, and deficiency-driven augmentation, our approach enables structured, scalable, and adaptive instruction data generation, laying the foundation for robust and generalizable instruction-tuned models.
@@ 6.2 Model Self-Improvement
Model self-improvement aims to iteratively enhance model capabilities through self-generated data or feedback signals. Representative approaches include self-refinement (madaan2023self; zhou2024cycle), multi-round generation and evaluation loops, and performance-driven data augmentation (yang2025knowing). These methods enable models to identify and address their own weaknesses by generating more challenging training samples, thereby facilitating continual improvement.
To ensure data quality, recent studies introduce deficiency diagnosis mechanisms (lighterness2024data), which analyze model performance on downstream tasks to detect knowledge gaps or skill deficiencies. These insights guide the targeted synthesis of training data(miller2025comparison). By incorporating such feedback into the training loop, models and datasets can co-evolve—improving both model performance and data coverage over time(bauer2024feedback).
Beyond simply increasing instruction quantity, recent research emphasizes instruction coverage (task/domain diversity) and depth (reasoning complexity). Several works have explored combining hierarchical tagging with complexity control to synthesize increasingly challenging and diverse instructions via evolutionary algorithms (lu2023instag; zhao2025beyond). These strategies enhance the cognitive depth of training data and promote models’ abilities in complex reasoning and compositional generalization.
@@ Appendix A Appendix
Given a dialogue between a user and a conversational assistant, and a classification standard that describes different abilities, determine which one or more major categories from this standard the assistant would need to utilize in order to complete the given dialogue. Note: only list the names of the abilities, and separate each different ability with angle brackets <><>. Classification Standard: standard